OpenStudy (phebe):

PEEEEEEEEEEEPSSSSSSSSSS!!!!!!!! HELPSSSSSSSSS Simplify by rationalizing the denominator: (-√18)/√12

OpenStudy (anonymous):

okay so you have \[\sqrt{18}\] you can write it as \[\sqrt{9*2}\]

OpenStudy (anonymous):

and you can write the denominator as \[\sqrt{4*3}\]

OpenStudy (phebe):

ok

OpenStudy (anonymous):

now you know that 9 is 3's squere so take 3 outside you get \[3\sqrt{2}\] and 4 is 2's squere so you get \[2\sqrt{3}\]

OpenStudy (phebe):

wait wats the answer

OpenStudy (anonymous):

now write these values in ur first number \[\frac{ -3\sqrt{2} }{ 2\sqrt{3} }\]

OpenStudy (anonymous):

its easy now simply solve it

OpenStudy (anonymous):

right?

OpenStudy (phebe):

thanx